HUBBLE SPACE TELESCOPE DAILY REPORT #2147 PERIOD COVERED: 0000Z (UTC) 06/17/98 - 0000Z (UTC) 06/18/98 Daily Status Report as of 169/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Four Sets of NIC/3 7919 (NICMOS Camera 3 Campaign Snapshots) The Near Infrared Camera (Camera Three) was used to make images of galaxies NGC-3271, NGC-5128, IC750, NGC-237 and NGC-4701 as part of the NICMOS Camera 3 campaign snapshots. The proposal is to obtain wide field snapshot images of nearby galaxies in Paschen Alpha which is inaccessible from the ground. The observations were executed as planned, and no anomalies were reported. 1.2 Completed WF/PC-2 7909 (POMS Test Proposal: WF II Parallel Archive Proposal) The WF/PC-2 was used to make parallel observations as part of the POMS test proposal. This test proposal is designed to simulate scientific plans. The observations were completed as planned, and no problems were reported. 1.3 Completed Two Sets of WF/PC-2 7618 (WF/PC-2 Cycle 7 Photometric Monitor UV/OPT STD) The WF/PC-2 was used to make observations of the standard star GRW+70D582 as part of a Cycle 7 photometric monitoring program. This ultraviolet spectrophotometric standard star was observed in a variety of filters and cameras to monitor the photometric stability and quantum efficiency of WF/PC-2 from the far-ultraviolet to near-infrared. The observations were completed as scheduled, and no problems were reported. 1.4 Completed STIS/CCD 7635 (STIS CCD Hot Pixel Annealing) The Space Telescope Imaging Spectrograph (CCD) was used to test the hot pixel annealing procedure. The effectiveness of the CCD hot pixel annealing process is assessed by measuring the dark current behavior before and after annealing and by searching for any window contamination effects. The observations were completed as planned, and no problems were reported. 1.5 Completed Three Sets of NIC/1/2/3 7596 (NICMOS MULTIACCUM Dark Characterization) The Near-Infrared Camera Multi-Object Spectrograph (Cameras One, Two and Three) was used to make a NICMOS MULTIACCUM dark characterization. The purpose of this proposal is to monitor the darks using three different MULTIACCUM sequences. The observations were executed as scheduled, and no anomalies were noted. 1.6 Completed Two Sets of NIC/1/2/3 7817 (A Complete NICMOS Map of the Hubble Deep Field) The Near Infrared Camera (Cameras One, Two and Three) was used to make observations to map the complete WF/PC-2 region of the central Hubble Deep Field. These data will provide unique information about the faintest objects, but will include only a small number of objects with spectroscopic redshifts or z ~ 3 Lyman break galaxies. During the second iteration of this proposal, a loss of lock at 168/0916Z, reported in HSTAR 6544 and in yesterday's report, occurred. Later in the same iteration there was an additional loss of lock, reported in HSTAR 6545 and 2.1 below. Otherwise, the observations were executed as scheduled, and no further anomalies were noted. 1.7 Completed Two Sets of WF/PC-2 8013 (WF/PC-2 Parallel Images to the HDF Flanking Fields {Companion Proposal to 7817}) The WF/PC-2 was used to make observations parallel to the Hubble Deep Field observations (i.e., flanking fields). The complete WF/PC-2 region of the central Hubble Deep Field {HDF} was being mapped with the NICMOS Camera 3 at F110W and F160W {proposal 7817}. While these observations were executing, the WF/PC-2 was working in parallel mode to image areas near this field. During the second iteration of this proposal, a loss of lock at 168/0916Z, reported in HSTAR 6544 and in yesterday's report, occurred. Later in the same iteration there was an additional loss of lock, reported in HSTAR 6545 and 2.1 below. Otherwise, the observations were completed as scheduled, and no further anomalies were reported. 1.8 Completed STIS/CCD 7926 (CCD Dark and Bias Monitor -- Continued) The Space Telescope Imaging Spectrograph (CCD) was used to make several dark and bias calibration observations. This is to measure CCD dark current and bias rates for the two supported amplifier settings {gain = 1 and gain = 4} in order to gather data for making superdarks and superbiases. These observations will also track the growth of hot pixels on the CCD in general and within the various target acquisition apertures specifically. The observations were executed as scheduled, and no anomalies were noted. 1.9 Completed STIS/CCD 7645 (STIS FUV-MAMA Cycle 7 Flats) The Space Telescope Imaging Spaectrograph (CCD) was used to make far-ultraviolet flat calibration images. This program will obtain FUV-MAMA flat-field observations with the Kr lamp for the construction of on-orbit D-flats for G140M. A dark calibration image was also taken. The observations were completed as planned, and no anomalies were reported. 1.10 Completed Four Sets of STIS/CCD 7908 (POMS Test Proposal: STIS Non-Scripted Parallel Proposal) The Space Telescope Imaging Spectrograph (CCD) was used to make calibration observations as part of the POMS Test Proposal. This was a STIS non-scripted parallel proposal. The observations were completed as planned, and no problems were reported. 1.11 Completed NIC/1/2/3 7474 (Searching for Low Mass Stars: the Mass Function at the Hydrogen Burning Limit) The Near Infrared Camera (Cameras One, Two, and Three) was used to use the unprecedented IR imaging capabilities of NICMOS to study the low-mass stars at the end of the main sequence in the nearby globular cluster NGC-6397 which spanning a wide range of metallicity, precisely at the wavelength corresponding to its maximum emission. These observations would complete our ongoing investigation by making it possible, for the first time, to reliably count stars right down to and beyond the main sequence H-burning limit, thus leaving no doubt as to whether the IMF indeed drops there or starts rising once again close to the limit. The observation completed as planned with no problems. 1.12 Completed NIC/2 7366 (Intense Galactic X-ray Sources in Crowded Fields) The Near-Infrared Camera (Camera Two) was used to make observations of the intense galactic x-ray, variable, binary, low mass star BR-Cir. This source lies in a severely optically crowded field, and there is strong evidence that past work has been deceived by spurious superpositions of unrelated objects. The observations were executed as scheduled, and no anomalies were noted. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 10 Successful : 10 Scheduled Reacquisitions : 8 Successful : 8 HSTAR 6544 and the previous report detailed a loss of lock at 168/0916Z. The proposals described in 1.6 and 1.7 above were affected. HSTAR 6545 doccuments a loss of lock during a profile (T3 slew) at 168/105259Z. The proposals described in 1.6 and 1.7 were affected. 2.2 FHST Updates: Scheduled : 19 Successful : 19 2.3 Operations Notes: Using ROP SR-1, the SSR EDEAC error counter was cleared two times. A real-time map was commanded at 168/1038Z and at 169/1956Z. Twenty-two minutes twenty-nine seconds of real-time engineering data was lost beginning at 168/164631Z due to operator error. This occurred from when the HST transitioned from mode 1 to mode 2 following the completion of a tracking/ranging event until engineering data recording was begun at 168/170900Z for a known ZOE. HSTAR 6546 was written. A series of five planned slews was uplinked from 168/1838Z until 168/1951Z to support current NICMOS observations. Per operations request 13830, a fix to the clearing slew group was uplinked at 168/1956Z. Per an operations request and ROP RD-7, ESTR reconditioning was performed at 169/0440Z. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH